# SplineSketch: Even More Accurate Quantiles with Error Guarantees

### v1

@@ 5. Experimental Evaluation
Figs. 1 and 3 show the average errors of the sketches depending on their size in bytes; see App. D for additional plots with average error (Fig. 8) and plots with the maximum error (Fig. 9). In all these plots, the error is presented in absolute terms, i.e., it is the absolute value of the difference between the true rank and the estimated rank of a query item, averaged or maximized over $10^{5}$ queries.
Overall, SplineSketch consistently provides the best accuracy if given sufficient size, namely $k\geq 100$, even though in rare cases or for small sketch sizes, $t$-digest or MomentSketch are better. In more detail, compared to $t$-digest, SplineSketch typically achieves 2-20 times smaller error, and in some cases, such as for normally distributed data, up to 100 times. On the other hand, there were cases in which $t$-digest performed better, such as for normal distribution followed by high-frequency items or in general, for datasets with skewed frequency distributions. In several other cases, their error was similar, e.g., the maximum error for signed loguniform distribution with large exponents. We note that the final $t$-digest size is somewhat unpredictable as the final number of centroids does not match the compression parameter given to it, sometimes by 66%. This demonstrates that $t$-digest does not use the space budget efficiently, unlike SplineSketch which always has $k$ buckets.
MomentSketch is mainly intended for use with very small size, typically up to $k=15$ moments and log-moments, when it often gives the best accuracy by 1-2 orders of magnitude. It works the best on many smooth distributions (uniform, normal, loguniform, etc.). However, its performace is significantly affected by non-smoothness as witnessed on the real-world datasets. Moreover, one cannot increase accuracy by increasing $k$ because of numeric issues. Furthermore, MomentSketch query procedure failed with an exception on many datasets, specifically on signed loguniform distribution with large exponents for any $k$ and on other datasets for large $k$. Finally, the KLL sketch is worse than SplineSketch typically by 1-3 orders of magnitude as it uses no interpolation (cf. (Schiefer et al., 2023) for KLL with interpolation; however, no implementation is available). Nevertheless, as KLL is comparison-based, KLL’s error depends just on the data size, not on their particular distribution, and decreases linearly with KLL’s size.
Figure 4. Log-log plot with time per update in $\mu$s depending on data size with fixed sketch size on normally distributed data.
Figure 5. Log-log plot with time per query in $\mu$s depending on the number of queries on $n=10^{8}$ samples from normal distribution.
@@ 6. Conclusions and Discussion
Besides very high accuracy on various datasets, the key advantage of our approach is flexibility. For instance, it is possible to resize the sketch, making it smaller or larger depending on how the memory resources change (e.g., one can use larger size $k$ for input processing than for storage). More importantly, if the user has a prior knowledge about the data distribution, it is possible to preset the initial thresholds of SplineSketch based on this prior distribution and aim for even lower error. Furthermore, if the prior knowledge turns out to be wrong, our sketch adapts to the new distribution, similarly as we have shown in Fig. 3(e).
One of the challenges when implementing SplineSketch is dealing with high-frequency items. We have tweaked the implementation to deal with them if they arrive at the beginning of processing a stream; however, if a high-frequency item starts occurring in the middle, it is unclear how to maintain the thresholds and counters in a right way. Still, if one processes inputs with such items, it is possible to deal with them by using a standard heavy-hitter algorithm, such as the MG sketch (Misra and Gries, 1982). Our implementation of SplineSketch is so far a prototype (in Python and Java) intended to demonstrate the advantages of our approach. Due to the intricacies of maintaining the buckets, our sketch is slower in processing the input stream than KLL or MomentSketch; however, a well-optimized Java version may be significantly faster than $t$-digest.
From the theoretical point of view, we ask whether one can modify our algorithm so that the error bounds have better dependency on $\alpha$. We note that there is some evidence from the comparison-based lower bounds in (Cormode and Veselý, 2020) that a factor of $\log\log\alpha$ may be necessary without packing more buckets into one memory word, similarly as done for q-digest in (Gupta et al., 2024). The accuracy in practice, despite being very high already, may also be improved, and it is not clear where is the accuracy limit of quantile summaries on a real-world datasets.
Our work opens up a new direction for future work in quantile streaming estimation. In particular, we believe that our techniques have a potential to yield much better quantile sketches with relative error guarantees, capturing the distribution tails more accurately than the median. However, achieving the relative error is substantially harder as witnessed by several major open problems in the theory of relative-error quantile sketches (cf. (Cormode and Veselý, 2020; Gribelyuk et al., 2025)).
@@ Acknowledgments.
A. Łukasiewicz and P. Veselý were partially supported by the ERC CZ project LL2406 of the Ministry of Education of Czech Republic. Part of this work was done while A. Łukasiewicz was a PhD student at the University of Wrocław. J. Tětek was supported by the VILLUM Foundation grant 16582. This research was partially funded from the Ministry of Education and Science of Bulgaria (support for INSAIT, part of the Bulgarian National Roadmap for Research Infrastructure). Part of this work was done when J. Tětek was visiting at the University of Wrocław and at Charles University. Part of this work was done while J. Tětek was employed at BARC, University of Copenhagen. P. Veselý was partially supported by Czech Science Foundation project 24-10306S and by Center for Foundations of Modern Computer Science (Charles Univ. project UNCE 24/SCI/008).
@@ Appendix A Bounds on Middle Point of Monotonic Cubic Polynomials
The following lemma provides a key technical property of PCHIP interpolation that we use subsequently in the analysis of our algorithm.
@@
Let $P(x)$ be a cubic polynomial such that $P(0)=0$, $P(1)=1$, and $P(x)$ is non-decreasing on the interval $[0,1]$. Then, $P\left(\dfrac{1}{2}\right)\in[1-\beta,\beta]\approx[0.066,0.933],$ where $\beta=\frac{1}{2}+\frac{\sqrt{3}}{4}\approx 0.933$.
@@
We prove that $P(\dfrac{1}{2})\leq\frac{1}{2}+\frac{\sqrt{3}}{4}$. The other inequality then holds by symmetry, that is, considering the polynomial $Q(x)=1-P(1-x)$ which is also non-decreasing and $Q(\dfrac{1}{2})=1-P(\dfrac{1}{2})$. From now on, we focus on proving that $P(\dfrac{1}{2})\leq\frac{1}{2}+\frac{\sqrt{3}}{4}$.
@@
First, using the constraint $a+b+c=1$, we express $c$ as $c=1-a-b$. Substituting this into the inequality $3a\lambda^{2}+2b\lambda+c\geq 0$, we obtain $3a\lambda^{2}+2b\lambda+(1-a-b)\geq 0$. Simplifying this expression gives:
@@
We now express $f(a,b,c)$ in terms of $a$ and $b$. Since $c=1-a-b$, we have:
@@
Multiplying both sides by $-\dfrac{3}{4}$ given:
$-\left(\dfrac{3a}{8}+\dfrac{b}{4}\right)\leq\dfrac{3}{4\sqrt{3}}.$
@@ Appendix B Proof of Uniform Error Guarantees and Full Mergeability
In the subsequent analysis, we assume that $k\geq 6$. The precise constant factors in the analysis depend on the interpolation used. Let $\beta>0$ be such that after splitting a bucket, the counters of both new buckets are smaller than the original counter by at least a factor of $1-\beta$. By Lemma 1, we have that $\beta=\frac{1}{2}-\frac{\sqrt{3}}{4}\approx 0.066$ for any monotonic cubic interpolation used to answer queries in the algorithm. For the linear interpolation we have $\beta=0.5$.
@@ B.1. Mergeability setting
For analyzing sketches created by a sequence of pairwise merge operations, it is convenient to define a binary tree $T_{0}$ with $n$ leaves, each representing one data item and inner nodes with two children corresponding to the merge operations. Additionally, we will use inner nodes representing the operations with buckets, namely splitting a bucket or joining a pair of adjacent buckets. That is, we define $T_{0}$ with root representing the final sketch, leaves corresponding the single data items, and inner nodes of three types: (1) merge-nodes with two children, representing merge operations such that the sketch represented by a merge-node is obtained by the union of buckets from the two source sketches represented by the children, as described in Section 3.2, and (2) split-nodes representing one split of a bucket, and (3) join-nodes representing one join of two adjacent buckets; split-nodes and join-nodes have a single child. Each node represents the sketch after the operation is performed. (A technicality is that the number of buckets may possibly be $k-1$ or $k+1$ after a join or split.) Other operations inside the sketch, namely creating initial buckets by equally spaced selection and merging buffer into buckets, need not be represented in $T_{0}$ for our analysis (while equally spaced selection affects the error, it is in total only by $n/k$).
@@ B.2. Existence of Joinable Pair of Buckets
In describing the consolidate subroutine (Section 3.1), we assumed that there is always a pair of buckets that we can join if a bucket counter exceeds the bound (1). Now we show that it holds.
@@
Consider any SplineSketch represented by a node $t$ of the merge tree $T_{0}$, and suppose that there is a splittable bucket $i$ according to Definition 2. Then there exists a joinable pair of adjacent buckets, according to Definition 1.
@@
First, observe that if bucket $i$ is splittable but does not exceed the bucket bound (1), there is a joinable pair of adjacent buckets by Definition 2. In the following, we thus assume that bucket $i$ exceeds the bound.
Note that a pair of adjacent buckets is not joinable for one of two reasons: (1) due to bucket bounds, namely that the resulting bucket would have counter exceeding $0.75\cdot C_{b}\cdot n/k$ or (2) due to protected thresholds. Let $n_{t}$ be the number of leaves in the subtree of $t$, i.e., the number of data items summarized by the sketch in $t$. Observe that there are at most $62n_{t}/(0.75\cdot C_{b}\cdot n_{t}/k)=\frac{8}{3}k/C_{b}\leq k/$ 6 pairs of buckets satisfying case (1), where we use $C_{b}\geq 16$. Thus, the main part of the proof is bounding the number of protected thresholds.
Let $j$ be the label of node $t$. By 1, $j$-nodes in the subtree of $t$ form a path $P_{j}$. If the sketches represented by nodes on $P_{j}$ all have at least $k/3+2$ joinable pairs of buckets, then the lemma holds. Consider the highest node $t^{\prime}$ on the path $P_{j}$ such that the sketch represented by $t^{\prime}$ has at least $k/3+2$ joinable pairs of buckets; we have that $t^{\prime}\neq t$, i.e., $t^{\prime}$ is below $t$, otherwise the lemma holds. Note that the sketch of the lowest node $u$ on $P_{j}$ has at most $k-1-k/6$ pairs that cannot be joined as there are no protected buckets, thus $t^{\prime}$ is well-defined since $k-1-k/6\geq k/3+2$ by $k\geq 6$.
@@
where $b_{i}$ is the counter of bucket $(\tau_{i-1},\tau_{i}]$ of the sketch of $v$. Clearly, $\Psi(v)\in[0,n_{v}]$ and moreover, the total increase of the potential on path $P^{\prime}_{j}$ is at most $n_{t}$. Next, observe that the potential does not increase by joining buckets (by Def. 1) and that every split at node $v$ of a bucket $i$ with counter $b_{i}>C_{b}\cdot n_{v}/k$ results in two buckets with counters at most $(1-\beta)\cdot b_{i}$. Since $\beta<0.25$, a split thus decreases the potential by at least
@@
We choose $C_{b}$ so that the number of split-nodes on path $P^{\prime}_{j}$ is at most $k/18$. As every split results in three protected thresholds, the number of thresholds that get protected on path $P^{\prime}_{j}$ is at most $k/6$. Since there are at least $k/3+2$ joinable pairs of buckets in node $t^{\prime}$, at most $k/6$ of these pair get non-joinable due to falling in case (1), and at most $k/6$ of them get non-joinable due to protected thresholds, there will be a joinable pair of buckets in node $t$. ∎

### v2

@@ 5.1. Evaluation in the Streaming Setting
We first present the results in the streaming setting, focusing on synthetic and real-world inputs without frequent items. Figs. 1 and 4 show the average errors of the sketches depending on their size in bytes. Fig. 5 shows the maximum error over all queries, again depending on the sketch size. In all these plots, the error is normalized by the input size, i.e., it is the difference between the true rank and the estimated rank of a query item divided by the input size $n$, averaged or maximized over $10^{5}$ queries. We also depict the theoretically optimal error bound of $O(1/k)$ in purple (loosely dotted), where $k$ is the sketch size and the hidden constant factor is based on the maximum error of KLL and GKAdaptive.
Overall, SplineSketch consistently provides the best accuracy if given sufficient size, namely $k\geq 100$, even though in rare cases or for small sketch sizes, $t$-digest or MomentSketch are better. In more detail, compared to $t$-digest, SplineSketch typically achieves 2-20 times smaller error, and in some cases, such as for normally distributed data, up to 100 times. In a few cases, their error was similar, e.g., the maximum error for signed loguniform distribution with large exponents. We note that the final $t$-digest size is somewhat unpredictable as the final number of centroids does not match the compression parameter given to it, sometimes by 66%. This demonstrates that $t$-digest does not use the space budget efficiently, unlike SplineSketch which always has $k$ buckets.
MomentSketch is mainly intended for use with very small size, typically up to $k=15$ moments and log-moments, when it often gives the best accuracy by 1-2 orders of magnitude. It works the best on many smooth distributions (uniform, normal, loguniform, etc.). However, its performance is significantly affected by non-smoothness as witnessed on the real-world datasets. Moreover, one cannot increase accuracy by increasing $k$ because of numeric issues. Furthermore, MomentSketch query procedure failed with an exception on many datasets, specifically on signed loguniform distribution with large exponents for any $k$ and on other datasets for large $k$. Finally, the KLL sketch and GKAdaptive are worse than SplineSketch typically by 1-3 orders of magnitude as they use no interpolation (cf. (Schiefer et al., 2023) for KLL with interpolation; however, no implementation is available). Nevertheless, as KLL and GKAdaptive are comparison-based, their errors depend just on the data size and the arrival order, not on their particular distribution, and their error decreases linearly.
Finally, we note that both versions of SplineSketch (without and with the MG sketch) provide the same accuracy on the inputs in Figs. 4 and 5.
@@
(a) Normal distribution
(b) Uniform distribution
(c) Pareto distribution
(d) Lognormal distribution
(e) Loguniform distribution
(f) Signed loguniform distribution with large exponents
(g) Normal distribution with a small parameter change after $n/2$ items
(h) Normal distribution with a large parameter change after $n/2$ items
Figure 4. Average rank error (log-scale) depending on the sketch size in bytes (average taken over $10^{5}$ evenly spaced queries).
(a) HEPMASS dataset (Whiteson, 2016)
(b) Power dataset (Hebrail and Berard, 2006)
(c) Books dataset (Kipf et al., 2019; Marcus et al., 2020)
(d) Normal distribution
(e) Uniform distribution
(f) Pareto distribution
(g) Lognormal distribution
(h) Loguniform distribution
(i) Signed loguniform distribution
(j) Signed loguniform distribution with large exponents
(k) Normal distribution with a small parameter change after $n/2$ items
(l) Normal distribution with a large parameter change after $n/2$ items
Figure 5. Maximum rank error (log-scale) depending on the sketch size in bytes, on a selection of datasets. The maximum is taken over $10^{5}$ evenly spaced queries.
[…]
(a) Log-log plot with time per query depending on the number of queries.
(b) Time per update depending on sketch size.
(c) Time per query depending on sketch size for $10^{5}$ queries.
[…]
5.2. Evaluation of Mergeability
We evaluated the performance sketches when they are created by pairwise merge operations. To this end, we split the input into 10 000 chucks, feed each chunk into one sketch, and then merge the resulting sketches into one in a binary tree fashion, i.e., merging sketches summarizing a similar number of items.
[…]
Figure 8. Average rank error (log-scale) depending on the sketch size in bytes, on the normal distribution for sketches built by merge operations from sketches for 10 000 input chunks.
Figure 9. Time per merge operation, in the same setting as in Figure 8.
5.3. Inputs with Heavy Hitters
Here, we focus on inputs with high-frequency items, where SplineSketch without the MG sketch does not perform so well. Our results in Figs. 10(a) and 10(b) confirm that using the Misra-Gries (MG) sketch alongside SplineSketch does indeed restore the very high accuracy when the MG sketch is large enough to detect the heavy hitters (note that the size of the MG sketch after compressing as described in Sec. 4.2 is accounted for). In more detail, while the error of $t$-digest and SplineSketch without MG gets worse and does not improve with the sketch size, SplineSketch with MG achieves by up to three orders of magnitude better error, and zero error on inputs consisting of a small number of distinct items only as in Fig. 10(b). We note that the error of KLL and GKAdaptive remains the same as on inputs with distinct items as they are comparison-based. As shown in Fig. 10(c), the update time of SplineSketch with MG is about three times worse than without MG (though it can be optimized), while merge and query times are only slightly larger.
(a) Average rank error (log-scale), on normal distribution with $n/2$ items, followed by $n/2$ high-frequency items.
(b) Average rank error (log-scale), on a sorted input with frequent items.
[…]
Figure 10. Accuracy and update time for inputs with frequent items, depending on the sketch size in bytes.
5.4. Ablation Studies
Here, we provide insights into the specific design choices in SplineSketch. First, we evaluated the benefit of using PCHIP interpolation and found that PCHIP interpolation achieves 3–10 times better error than just linear interpolation over the buckets; see Fig. 11(a).
Second, we focused on the heuristic error and compared several variants instead of the second derivative of the empirical CDF, normalized by length squared, defined in (2). The evaluated alternatives are: using no heuristic error, bucket length, bucket counter (which is the first derivative of the empirical CDF), and the third derivative of the empirical CDF, normalized by length cubed. Fig. 11(b) shows that (2) is indeed the best, while using the bucket length, i.e., making the buckets similar in length on the input range, is only about 20% worse. The third derivative is about 1.5–2 times worse, and using the bucket counter or no heuristic error is over 10 times worse.
Finally, we briefly comment on parameter tuning. We did not observe significant differences in the accuracy of SplineSketch with respect to parameter $\gamma>1$ in Definition 2 (for values between 1.1 and 2). For the initial multiplicative constant $C_{b}$ in the bucket bound (1), we tested values in $[1,7]$ and for $C_{b}\geq 3$, the accuracy was similar ($C_{b}<3$ was only better on the uniform distribution). Other parameters, such as the epoch increase factor (default=1.25) or the minimum fraction of the bucket bound (default=0.01) also do not influence the accuracy much if adjusted. See the supplementary repository for details.
(a) Interpolation.
(b) Heuristic error.
Figure 11. Ablation studies on the normal distribution.
@@ 6. Conclusions and Discussion
Besides very high accuracy on various datasets, the key advantage of our approach is flexibility. For instance, it is possible to resize the sketch, making it smaller or larger depending on how the memory resources change (e.g., one can use larger size $k$ for input processing than for storage). More importantly, if the user has a prior knowledge about the data distribution, it is possible to preset the initial thresholds of SplineSketch based on this prior distribution and aim for even lower error. Furthermore, if the prior knowledge turns out to be wrong, our sketch adapts to the new distribution, similarly as we have shown in Fig. 4(h).
One of the challenges when implementing SplineSketch is dealing with high-frequency items. One possible approach is using the Misra-Gries sketch (Misra and Gries, 1982), which restores theoretical guarantees (Sec. 3.3) and very high accuracy for inputs with heavy hitters, as demonstrated in Section 5.3. Nevertheless, using a heavy-hitter sketch comes at a cost of increased update, merge, and query times, so we also provide a faster version without a heavy hitter sketch. Another limitation of our implementation arises from using the double data type in Java, which means that for some inputs with many significant digits (e.g., the SOSD benchmark (Kipf et al., 2019; Marcus et al., 2020)), accuracy degrades as the sketch rounds many items to the same double value; this can be addressed by changing the underlying data type, e.g., to 64-bit integers. While we optimized the Java version to some extent, our implementation is so far a prototype intended to demonstrate the main advantages and challenges of our approach. Due to the intricacies of maintaining the buckets, our sketch is slower in processing the input stream than KLL or MomentSketch but already faster in streaming than $t$-digest. However, further optimizations are possible, such as the ones presented in (Pfeil et al., 2025) (though not applicable to Python or Java). Thus, a well-optimized implementation may be significantly faster than $t$-digest.
From the theoretical point of view, we ask whether one can modify our algorithm so that the error bounds have better dependency on $\alpha$. We note that there is some evidence from the comparison-based lower bounds in (Cormode and Veselý, 2020) that a factor of $\log\log\alpha$ may be necessary without using bit-packing tricks developed for $q$-digest in (Gupta et al., 2024). The accuracy in practice, despite being already very high, may also be improved, and it is not clear where is the accuracy limit of quantile summaries on real-world datasets.
Our work opens up a new direction for future work in quantile streaming estimation. In particular, we believe that our techniques have potential to yield much better quantile sketches with relative-error guarantees, capturing the distribution tails more accurately than the median. However, achieving the relative error is substantially harder as witnessed by several major open problems in the theory of relative-error quantile sketches (cf. (Cormode and Veselý, 2020; Gribelyuk et al., 2025)).
@@ Acknowledgments.
We thank the reviewers for their constructive feedback and suggestions for improvement. ChatGPT was used to generate parts of this work, including polishing language in the text, pseudocodes prototyping, and translation of SplineSketch from Python to Java. We have checked all the LLM generated content, corrected it where necessary. In particular, the Java implementation was largely rewritten and optimized manually.
@@ Appendix A Bounds on Middle Point of Monotonic Cubic Polynomials
The following lemma provides a key technical property of PCHIP interpolation that we use subsequently in the analysis of our algorithm in Appendix B.
@@
Let $P(x)$ be a cubic polynomial such that $P(0)=0$, $P(1)=1$, and $P(x)$ is non-decreasing on the interval $[0,1]$. Then, $P\left(\dfrac{1}{2}\right)\in[\beta,1-\beta]\approx[0.066,0.933],$ where $\beta=\frac{1}{2}-\frac{\sqrt{3}}{4}\approx 0.067$.
@@
We prove that $P(\dfrac{1}{2})\leq\frac{1}{2}+\frac{\sqrt{3}}{4}$. The other inequality then holds by symmetry, that is, considering the polynomial Q(x)=1P(1x)Q(x)=1-P(1-x) which is also non-decreasing and $Q(\dfrac{1}{2})=1-P(\dfrac{1}{2})$. From now on, we focus on proving that $P(\dfrac{1}{2})\leq\frac{1}{2}+\frac{\sqrt{3}}{4}$.
@@
First, using the constraint $a+b+c=1$, we express $c$ as c=1abc=1-a-b. Substituting this into the inequality $3a\lambda^{2}+2b\lambda+c\geq 0$, we obtain $3a\lambda^{2}+2b\lambda+(1-a-b)\geq 0$. Simplifying this expression gives:
@@
We now express $f(a,b,c)$ in terms of $a$ and $b$. Since c=1abc=1-a-b, we have:
@@
Multiplying both sides by $\dfrac{3}{4}$ given:
$\left(\dfrac{3a}{8}+\dfrac{b}{4}\right)\leq\dfrac{3}{4\sqrt{3}}.$
@@ Appendix B Proof of Uniform Error Guarantees and Full Mergeability
In the subsequent analysis, we assume that $k\geq 6$. The precise constant factors in the analysis depend on the interpolation used. Let $\beta>0$ be such that after splitting a bucket, the counters of both new buckets are smaller than the original counter by at least a factor of $1-\beta$. By Lemma 1, we have that $\beta=\frac{1}{2}-\frac{\sqrt{3}}{4}\approx 0.067$ for any monotonic cubic interpolation used to answer queries in the algorithm. For the linear interpolation we have $\beta=0.5$.
@@ B.1. Mergeability setting
For analyzing sketches created by a sequence of pairwise merge operations, it is convenient to define a binary tree $T_{0}$ with $n$ leaves, each representing one data item and inner nodes with two children corresponding to the merge operations. Additionally, we will use inner nodes representing the operations with buckets, namely splitting a bucket or joining a pair of adjacent buckets. That is, we define $T_{0}$ with root representing the final sketch, leaves corresponding the single data items, and inner nodes of three types: (1) merge-nodes with two children, representing merge operations such that the sketch represented by a merge-node is obtained by the union of buckets from the two source sketches represented by the children, as described in Section 3.2, and (2) split-nodes representing one split of a bucket, and (3) join-nodes representing one join of two adjacent buckets; split-nodes and join-nodes have a single child. Each node represents the sketch after the operation is performed. (A technicality is that the number of buckets may possibly be k1k-1 or $k+1$ after a join or split.) Other operations inside the sketch, namely creating initial buckets by equally spaced selection and merging buffer into buckets, need not be represented in $T_{0}$ for our analysis (while equally spaced selection affects the error, it is in total only by $n/k$).
@@ B.2. Existence of Joinable Pair of Buckets
In describing the consolidate subroutine (Sec. 3.1), we assumed that there is always a pair of buckets that we can join if a bucket counter exceeds the bound (1) or if there is a new minimum or maximum in the buffer. Now we show that it holds.
@@
For any node $t$ of the merge tree $T_{0}$, the SplineSketch represented by $t$ contains a joinable pair of adjacent buckets, according to Definition 1.
@@
Note that a pair of adjacent buckets is not joinable for one of two reasons: (1) due to bucket bounds, namely that the resulting bucket would have counter exceeding $0.75\cdot C_{b}\cdot n/k$ or (2) due to protected thresholds. Let $n_{t}$ be the number of leaves in the subtree of $t$, i.e., the number of data items summarized by the sketch in $t$. Observe that there are at most $2n_{t}/(0.75\cdot C_{b}\cdot n_{t}/k)=\frac{8}{3}k/C_{b}\leq k/6$ pairs of buckets satisfying case (1), where we use $C_{b}\geq 16$. Thus, the main part of the proof is bounding the number of protected thresholds.
Let $j$ be the label of node $t$. By 1, $j$-nodes in the subtree of $t$ form a path $P_{j}$. If the sketches represented by nodes on $P_{j}$ all have at least $k/3+2$ joinable pairs of buckets, then the lemma holds. Consider the highest node $t^{\prime}$ on the path $P_{j}$ such that the sketch represented by $t^{\prime}$ has at least $k/3+2$ joinable pairs of buckets; we have that $t^{\prime}\neq t$, i.e., $t^{\prime}$ is below $t$, otherwise the lemma holds. Note that the sketch of the lowest node $u$ on $P_{j}$ has at most k1k/6k-1-k/6 pairs that cannot be joined as there are no protected buckets, thus $t^{\prime}$ is well-defined since $k-1-k/6\geq k/3+2$ by $k\geq 6$.
@@
where $b_{i}$ is the counter of bucket $(\tau_{i-1},\tau_{i}]$ of the sketch of $v$. Clearly, $\Psi(v)\in[0,n_{v}]$ and moreover, the total increase of the potential on path $P^{\prime}_{j}$ is at most $n_{t}$. Next, observe that the potential does not increase by joining buckets (by Def. 1) and that every split at node $v$ of a bucket $i$ with counter $b_{i}>C_{b}\cdot n_{v}/k$ results in two buckets with counters at most $(1-\beta)\cdot b_{i}$. Since $\beta<0.25$ [^1] In case $\beta\geq 0.25$, the constant factors in this proof need to be adjusted., a split thus decreases the potential by at least
@@
We choose $C_{b}$ based on $\beta$ so that the number of split-nodes on path $P^{\prime}_{j}$ is at most $k/18$. As every split results in three protected thresholds, the number of thresholds that get protected on path $P^{\prime}_{j}$ is at most $k/6$. Since there are at least $k/3+2$ joinable pairs of buckets in node $t^{\prime}$, at most $k/6$ of these pair get non-joinable due to falling in case (1), and at most $k/6$ of them get non-joinable due to protected thresholds, there will be a joinable pair of buckets in node $t$. ∎